OpenStudy (anonymous):

I need help with this: m-13=-41.

OpenStudy (amistre64):

what do you think the first step in solving this would be?

OpenStudy (anonymous):

you should solve for "m", so you'll have to move what? ^_^

OpenStudy (anonymous):

just move the -13 to the other side, making it positive, so it will look like this m = -41+13

OpenStudy (anonymous):

and the answer is m = -28

OpenStudy (amistre64):

understand that a number add to its own opposite value equals zero: 3+(-3) = 0 -5 + (5) = 0 -13+(13) = 0

OpenStudy (anonymous):

thank you angoo and amistre

OpenStudy (amistre64):

yw .... i think :)

OpenStudy (amistre64):

my daughter hates for me to help her with her math homework; she would rather I DO the work for her and give her an answer. Instead I let her come to an answer with a little guidance.... just a thought..

OpenStudy (anonymous):

lol, never do that!